March BG Com Night 1, March 6-7 2009 MST

Observers: J. Hill (Tucson), A. Rakich (mtn)
Lucifer support: M. Juette
Telescope Operator: W. Wack
SW Support: N. Cushing, C. Biddick, M. de la Pena, P. Grenz, T. Leibold, S. Hooper (Tucson) plus J. Borelli and M. Jutte (Germany)
Telescope Support: J. Rousseau

Summary:

Closed for weather, but we are still able to do many aspects of software testing.

The major unresolved problem is that the M3 UMAC selector axis is faulting. We need to get this working, or at least in correct position to get a star image.

Michele is fighting with a problem in the back-calculated RA/DEC in the new PCS. The values flip between the correct ones and NaNs. At the moment it appears to be a problem related to the rotator coordinates coming from MCSPU.

Details (Times reported in MST)

Preparations

19:00 Andrew has finished his on-site AGW testing. Weather is bad with clouds and high winds, so Andrew is returning to Tucson. D. Ashby is balancing the telescope with the AGW access platform still in place.

TCS testing - Build 22

20:26 IIFGUI - PCS is reporting alternating good and bogus RA/DEC coordinates.

M3 UMAC Issues

20:28 Paul is working on SX M3 UMAC.

20:30 pointto BS9108 TRACK - preset failed because M3 "not at a focal station"

20:35 Paul sends Wayne up to reset M3 UMAC. Third LED was blinking 3 times.

20:49 Paul sends Wayne up to reset M3 UMAC again. Fourth LED was blinking 9 times.

20:50 Michele is deploying an older patch version of PCS to see if it cures the NaN garbage problem in the RA/DEC back calculation. It did, so now she is going to try to fix the new version of PCS.

21:03 Selector amp faults again. Paul sends Wayne up to reset M3 UMAC again.

I gave up counting after this, but they cycled power at least 10 times.

22:00 We have tested that irc commands to IIF are working through the new ICE interface of IIF.

23:00 Tim and Wayne replaced the failed 24V fan on the M3 UMAC with two 12V fans. Unfortunately, that has not helped the problems of the selector amp faults.

23:14 pointto BS9126 TRACK - failed

GCSL is in star image simulation mode. M3 is not in position.

23:19 Change lbt.conf so SX M3 is in simulation mode.

23:24 pointto BS9126 TRACK - failed, IIF still thinks M3 is not at a focal station. Paul is working on improvements to M3 simulation mode.

23:40 We moved Polycom ....27 back from conference room N505 to the SW TO-station so we can talk to Jose and Marcus at the same time. We were careful to switch to the alternate power connection.

Unfortunately, Jose's typical connection time is only 10 sec.

00:00 Open vent doors and turn on mirror ventilation.

00:10 pointto BS9126 TRACK - succeeded with the new version of OSS.

GCS tests in starOverlay mode

00:12 pointo BS9126 gs=0 GUIDE - succeeded with simulated seeing of 0.19 arcsec.

00:17 pointto BS9126 gs=0 ACTIVE - succeeded

IIF problems when LUCIFER connects

00:22 IIF dies for no apparent reason, but observing continued! (It was Marcus.)

00:25 restart IIF

Neither Marcus nor Jose are able to connect for more than 10 sec on the Polycom. Not sure if this is an Arizona problem or something in the trans-atlantic connection.

00:31 IIF died again - this was when Marcus tried to connect with LUCIFER software.

01:23 Jose and Marcus sorted out the IIF problem. It had to do with doing a getMultiparam without being authorized and getting a null pointer. For now, LUCIFER needs to authorize before asking for multiparams. (This also means that we can't use the DIMM right now.) Jose will work on it more tomorrow after he has had some sleep.

Left primary mirror is routinely failing to collimate (I think it is chase not making tolerance.)

LUCIFER Preset Tests

01:35 Marcus sends a preset and it succeeds, but SX primary mirror panic'd because of too large correction forces. GCSL in simulation mode sends garbage Zernikes.

01:46 Another LUCIFER preset fails because Michele killed PCS.

TCS Test #5 mixed with LUCIFER tests.

02:04 Small offsets are taking 17-19 seconds. A fair fraction of that is checking the stability of the guide image. I think it did not update the primary for repeated small offsets, although the time consumed was the same.

02:10 We finish with Marcus for today after several successful presets and offsets. He was alos able to see the 0.1 arcsec offsets I was sending from IRAF.

02:15 Chiller LWT is still at -22.3 after ventilating both mirrors for more than two hours. It is very slowly losing ground, but only very slowly.

TCS Test #8

02:19 pointto BS9145 ACTIVE

02:21 elslewtohold 0

02:27 pointto BS9152 ACTIVE - succeeded without any goofy offloads being added.

02:29 elslewtohold 0 - EL axis stopped near horizon.

02:34 Close vent doors and stop mirror ventilation as humidity is rising.

02:39 pointto BS9145 GUIDE

02:43 offset 3 3 DETXY, but we are running PCS without the feature for Test 11.

03:00 Michele installs ECS patch to get the correct dewpoint readings.

03:15 pointto BS9145 GUIDE gs=0 - after Michele restarted PCS

03:16 pointto BS9145 GUIDE gs=2 - couldn't move probe there

03:18 pointto BS9145 GUIDE gs=4 - succeeded




http://wiki.lbto.arizona.edu/twiki/bin/view/TCS/TCSRemainingTests


-- JoarBrynnel - 02 Mar 2009
Topic revision: r6 - 16 Mar 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback